OpenStudy (clara1223):

simplify: sec(x)-tan(x)sin(x)+(sin(2x))/(2sin(x))

OpenStudy (clara1223):

\[\sec \theta-\tan \theta \sin+\frac{ \sin(2\theta) }{ 2\sin(\theta) }\]

OpenStudy (mertsj):

\[\frac{1}{\cos \theta}-\frac{\sin \theta}{\cos \theta}\times \sin \theta+\frac{2\sin \theta \cos \theta}{2\sin \theta}=\]

OpenStudy (clara1223):

answer choices are: a)cos(x)^2 b)2cos(x) c)sin(x)+1 d)cos(x)+1 e)2sin(x)

OpenStudy (mertsj):

\[\[\frac{1-\sin ^2\theta}{\cos \theta}+\cos \theta=\frac{\cos ^2\theta}{\cos \theta}+\cos \theta =\cos \theta +\cos \theta =\]

OpenStudy (mertsj):

\[2 \cos \theta\]

OpenStudy (clara1223):

@Mertsj Thanks so much!

OpenStudy (mertsj):

yw